The volume of a cube is given by V = s3, where s is the length of a side. Find the length of a side of a cube if the volume is 500 cm3. Round answer to three decimal places.
Answer by edjones(8007)

V=s^3=500
s=500^1/3 (cube root of 500)
s=7.937
Ed